OpenStudy (anonymous):

Evaluate the Riemann sum for (x) = x3 − 6x, for 0 ≤ x ≤ 3 with six subintervals, taking the sample points, xi, to be the right endpoint of each interval. Give three decimal places in your answer.

OpenStudy (freckles):

divide 3-0 by 6 to figure out what the length of each sub-interval should be and just take the right endpoint of each sub-interval...

OpenStudy (freckles):

then once you have the right endpoint of each sub-interval you can find the area of your 6 rectangles and add them up together

OpenStudy (anonymous):

So basically I have an interval every half?

OpenStudy (freckles):

yes example interval 1 would be [0,1/2] example 2 interval 2 would be [1/2,1] and so on... the right endpoints are 1/2,1,...

OpenStudy (freckles):

each sub-interval has length 1/2

OpenStudy (freckles):

so what would be the 3rd sub-interval?

OpenStudy (anonymous):

1.5 no?

OpenStudy (freckles):

do you mean [1,1.5] because a number by itself is not an interval

OpenStudy (freckles):

and the right endpoint of that interval is 1.5

OpenStudy (anonymous):

Yeah sorry.

OpenStudy (freckles):

[0,1/2] [1/2,1] [1,3/2] [3/2,2] [2,5/2] [5/2,3] these are your 6 subinterval

OpenStudy (freckles):

you will consider the y values for each right endpoint of each of the sub-intervals

OpenStudy (freckles):

base of each rectangle is 1/2 you have to evaluate: \[\frac{1}{2}(f(\frac{1}{2})+f(1)+f(\frac{3}{2})+f(2)+f(\frac{5}{2})+f(3)]\]

OpenStudy (freckles):

I believed when you typed this "(x) = x3 − 6x" you meant f(x)=x^3-6x

OpenStudy (anonymous):

You are right. Let me evaluate that.

OpenStudy (anonymous):

Am I supposed to get a negative number?

OpenStudy (freckles):

that is what I have

OpenStudy (freckles):

we could get a better idea of why that is by graphing the thingy on [0,3]

OpenStudy (anonymous):

Ok, but it is equal to -3.94?

OpenStudy (freckles):

 it is definitely a net area that area below the x-axis must have been bigger then the area of above the x-axis since we got a negative result and yes -3.94 is what I have too